### v1

# Natasha Kirova    Serguei Brazovskii

Abstract
Phase transformations induced by short optical pulses is a mainstream in studies of of dynamics of cooperative electronic states. We present a semi-phenomenological modeling of spacio-temporal effects expected when optical excitons are intricate with the order parameter as it happens e.g. in organic compounds with neutral-ionic ferroelectric phase transitions. A conceptual complication appears here, where both the excitation and the ground sate ordering are built from the intermolecular electronic transfer. To describe both thermodynamic and dynamic effects on the same root, we adopt for the phase transition a view of the Excitonic Insulator - a hypothetical phase of a semiconductor which appears if the exciton energy becomes negative. After the initial pumping pulse, a quasi-condensate of excitons can appear as a macroscopic quantum state which then evolves interacting with other degrees of freedom prone to an instability. The self-trapping of excitons enhances their density which can locally surpass a critical value to trigger the phase transformation. The system is stratified in domains which evolve through dynamical phase transitions and may persist even after the initiating excitons have recombined.
1 Introduction
There is a persistent interest in phase transformations induced by optical pulses in cooperative electronic states (see e.g. a collection of articles [1], proceedings [2, 3] and the information from the latest conference [4]). The traditional technique of pump-induced phase transitions (PIPT) is based on the direct pumping between filled and empty electronic bands (the photon absorption energy $E_{ph}$ is above their separating gap $E_{g}$), creating the $e-h$ plasma to affect the correlated electronic states such as charge density waves (CDW), superconductors, or Peierls, Mott and excitonic insulators. The femtosecond shortness of pumps allows to resolve still disentangled ”fast” (10-100 fs) electronic and ”slow” (ps) lattice degrees of freedom. The intensity of pumps permits to kick the system far from the equilibrium which allows to follow the dynamical evolution over a wide range of the known phase diagram, see e.g. [5]. It becomes possible to reach also short-leaving ”hidden states” unavailable in thermodynamic approaches and even to obtain a truly stable one, see [6, 7].
The exploitation of generically insulating systems, with their neutral excitations - the excitons, is a still ill explored land. The field is unlimited since many insulators are close to electronic and/or structural or magnetic instabilities, and all non-metallic systems possess some excitons accessible for pumping. The sufficiently dense gas of such excitons may be ready to form a coherent quantum state close to the Bose–Einstein condensation (BEC) [8]. Thus, the PIPT science for insulators needs also to include studies of coherent effects in ensembles of excitons. The BEC of excitons was searched for in bulk semiconductors, particularly in $Cu_{2}O$ [9] and finally observed in bilayer heterostructures. The BEC of excitons in the bulk is restricted to a low temperature $T_{BEC}$ because of their typically low concentration $n_{ex}$ limited by a ”short” $\backsim 10^{-9}$sec. life time. This is why the BEC was realized only for much lighter particles - the polaritons which entangle the exciton and the photon (see [10] for a review). A successful escape from this limitation has come with design of artificial bilayer heterostructures where electrons and holes are space separated, see [11] or e.g. [12]. We warn that PIPT offer another option: in conditions of the femtosecond pumping and the picoseconds’ scale of observation, $n_{ex}$ grows to huge levels of about one excitation for a few unit cells, yielding expectations for $T_{BEC}$ up to $100K$. The ”short” life time becomes ”long” more than enough.
With a good luck, the pumping can go directly to the excitonic level by a resonant photon absorption $E_{ph}\approx E_{ex}$; then the excitations appear initially as delocalized plane waves, since photons create them at zero momentum $k=0$. In more common cases of a deviation from the resonance $E_{ph}>E_{ex}$, the hot excitons appear, still as delocalized waves, but with a distribution of momenta $k$, so that the phonon-assistant and/or collision processes are required for the energy relaxation. Cooling below the bandwidth of the excitonic spectrum $E_{ex}(k)$ towards its bottom (supposed to be at $k=0$) should lead to development of a quasi-condensate - the distribution peak at the lowest energy. The development of the BEC from the initially incoherent state is an intricate process, nowadays observed in cold atoms with many attempts to build a theory at both phenomenological [13] and microscopic [14, 15, 16] levels. The picture is not universal, depending on interactions of bosons (the pure repulsion for dipole excitons in double-wall structures, or the common, for excitons in the bulk, crossover among the attraction at large distances and the repulsion at short ones), and on inelastic scattering mechanisms required for cooling.
Almost at the same time as Keldysh suggested the idea about the BEC of excitons, the concept of an ”excitonic insulator” (EI) as a special phase of a semiconductor was introduced [17, 18] after a vague proposal in [19] and its first theoretical development in [20]. The EI became a standard name for a state formed by the condensate of bound electron–hole pairs, on top of a semiconducting or a semimetallic state (see basic reviews [21, 22] and more modern ones in the collection [23]). The state of the EI appears when the total energy of the exciton vanishes, $E_{ex}=E_{g}-E_{b}\rightarrow 0$. That implies the possibility of independent manipulation of the conduction gap $E_{g}$ and the exciton’s binding energy $E_{b}$ (e.g., by the composition or the pressure). There is some confusion on the terminology since recently the name EI was assigned to states of a kind of CDWs [24, 25, 26, 27] where the lattice component is not seen, so the CDW is due to exclusively to $2P_{F}$ modulations of the electronic density; there can be no excitons with the well screened Coulomb interactions.
Already in static conditions, the microscopic theory of the thermodynamic EI state and the theory of the BEC for optically pumped excitons are closely related, differing mostly by the monitoring parameters: the chemical potential $\mu_{ex}$ and the density $n_{ex}$ of excitons correspondingly. For the continuous optical pumping, this duality was strongly emphasized later [28].
A dynamical theory of PIPTs is challenging for ab initio microscopic approaches, e.g. [29, 30, 31]. Nevertheless, the evolution at longer time scales is controlled by collective variables, such as the order parameter, excitons’ concentration and lattice displacements. Then a phenomenological approach grant us the possibility to describe oscillatory and inhomogeneous regimes of a dynamical symmetry breaking, already proved to be in a good accordance with experiment (see e.g. [5]). Another example was the modeling [32] of a stable hidden state of a polaronic Mott insulator observed in $1T-TaS_{2}$ [7]. The phenomenological approach becomes inevitable when we consider time dependent spacially inhomogeneous regimes which finally emerge in modeling.
In this article we study the interplay among the pumping to excitons and the developing thermodynamic instabilities resulting in an induced phase transition. We present the phenomenological model which is based on a generalized concept of the EI formation evolving gradually from the BEC of excitons. On this bases we perform the modeling of spacio–temporal effects expected for the optically pumped ensemble of excitons coupled to the order parameter and lattice displacements. We shall briefly consider the case of extrinsic excitons which nature is essentially different with respect to the thermodynamical order parameter; the presence of excitons just shifts the transition temperature and the equilibrium value of the order parameters. Mainly we shall be concerned with a conceptually interesting case of intrinsic excitons which density contributes directly to the thermodynamic order parameter; this is the case of the charge transfer excitons with respect to the charge ordering transition. Keeping in mind the already existing experiments involving neutron-ionic transitions, we shall assume the second order character of the thermodynamic transition and include also possible interactions with another degree of freedom - the lattice displacements.
We shall report two major observations. 1. The interaction of excitons with the order parameter can lead to their self-trapping enhances their local density which can locally surpass a critical value to trigger the phase transformation. The system is stratified in domains which evolve through dynamic phase transitions and may persist even after recombination of pumped excitons. 2. The excitons (unlike well studied cold Bose atoms) allow for virtual processes of creation/annihilation of their pairs from/to the vacuum which lead to a dynamical transition of the phase locking, thus indicating the demarcation between regimes of BEC and the EI. 3. We shall model some effects of transformations among the condensed and incoherent reservoirs of excitons. First, we shall study stochastic equations emulating a random falling of hot particles to the condensate. Second, we shall model a two-fluid macroscopic coexistence with a gradual absorption of the initially pumped particles to the condensate. Preliminary results and discussions can be found in earlier authors’ publications [33, 34, 35, 36].
The article is organized as follows: Section 2 is alone devoted to case when the excitons are foreign with the order parameter. In Section 3 we the intricate case of the EI and the BEC. In Section 4 we present multi-field model for EI complemented by lattice displacements. Section 5 is dedicated to the effects coming from initially present hot excitons. The Section 6 is devoted to conclusions.
2 Extrinsic excitons and an arbitrary order parameter.
The conceptually simple and expectedly most common situation refers to the case of extrinsic excitons which nature is different with respect to that of the order parameters $\eta$. Their interaction just affects the transition temperature and the equilibrium value of $\eta$ while variations of $\eta$ shift the exciton energy.
Here and throughout the article, our basic assumption is that after a high initial pumping the quasi-condensate of optically pumped excitons appears sufficiently early as the macroscopic quantum state. The exciton is a quantum state in both the internal and the center of mass coordinates The condensate can be described by a macroscopic wave function $\Psi=|\Psi|\exp(i\zeta)$ which evolves interacting with other degrees of freedom prone to instability. The attempt to describe a coexistence of the BEC of excitons and their initially created incoherent liquid will be given below in Ch.5.
Suggesting for simplicity the smallness of $\eta$ in a proximity to the (nearly) second order phase transition from the high temperature (T) state with $\eta=0$, we can write the energy functional density for $\eta$ and $\Psi$ as
$W(\eta,\Psi)$ $=W(\eta)+(E_{ex}^{0}+g_{1}\eta+g_{2}\frac{\eta^{2}}{2})|\Psi|^{2}+\frac{k}{2}|\Psi|^{4}+\frac{\hbar^{2}}{2M}\left(\nabla\Psi\right)^{2}$ (1)
$W(\eta)$ $=\frac{a}{2}\eta^{2}+\frac{g_{3}}{3}\eta^{3}+b\eta^{4}+\frac{c}{2}\left(\nabla\eta\right)^{2}$ (2)
Here the odd in $\eta$ terms (the ones with coefficients $g_{1,3}$) are present if the order parameter is not symmetry breaking which is commonly a case of the Ist order transition close to the IInd order one (the critical point of monitoring parameters where $a=0$). The odd $\varpropto g_{3}\eta^{3}$ term with $g_{3}<0$ can give rise to the the second, above the first one at $\eta_{1}=0$, minimum at some $\eta_{2}>0$ providing the appearance of the metastable state which further on becomes the preferable one. At presence of excitons with a density $n_{ex}=|\Psi|^{2}$, their interaction with $\eta$ shifts the critical point from $a_{cr}=0$ to $a_{cr}^{{}^{\prime}}=-g_{2}n_{ex}$. Reciprocally, the appearance of $\eta\neq 0$ shifts the exciton energy from $E_{ex}^{0}$ to $E_{ex}^{0}+g_{2}\eta^{2}/2$. The most common situation in PIPT, at least with conventional mechanisms of pumping to e-h excitations, is that the pumping destroys temporarily the low temperature (T) phase which corresponds to increase of $a$ from the negative value $a<0$ to the positive one $a^{\prime}=a+g_{2}n_{ex}>0$, hence expecting $g_{2}>0$. That can be verified by observing the increasing value of the exciton level in the ordered state $\eta\neq 0$. This situation is not a dogma and the opposite sign $g_{2}<0$ may take place. Intuitively, the sign of $g_{2}$ is expected to be positive if $\eta$ drives the system towards further dielectrization and negative if the system is driven towards metallization.
Within PIPT preferences, the most interesting scenario is a reaching of the second (meta)stable state at $\eta_{2}\neq 0$ which is imperative for a subset of non-symmetry-breaking transitions. That expectation does not exclude completely some rare symmetry breaking case where the negative higher order therms, like the change of sign in the coefficient $k$, are known to occur. In these cases, before pumping or at the low one, the phase with $\eta=0$ is stable and the phase with $\eta=\eta_{2}$ is unstable. As the result of pumping, the barrier and both energy minima diminish their values shifting at the same time their positions to higher $\eta$.
The dynamics of this system can be studied by the set of equations
$i\hbar\partial_{t}\Psi+i\hbar\Gamma\Psi$ $=\frac{\delta W(\eta,\Psi)}{\delta\Psi^{\ast}}$ (3)
$\frac{a}{\omega^{2}}\frac{\partial^{2}\eta}{\partial t^{2}}+\gamma\frac{\partial\eta}{\partial t}$ $=-\frac{\delta W(\eta,\Psi)}{\delta\eta}$ (4)
[…]
Here $W$ is given by (1), $\omega$ is the bare frequency of oscillations of a weakly perturbed field $\eta$, $\gamma$ is their attenuation parameter, $\Gamma$ is the decay rate in the ensemble of excitons, $V$ is the system volume. Above, we have ignored effects breaking the conservation of the total number $N$ of excitons which will be important later for the case of intrinsic excitons, in Section 3.
The first critical pumping $n_{ex}=n_{c}$ corresponds to the binodal formation: the energies of both energy minima are equal, being still separated by the barrier. The second critical pumping at $n_{ex}=n_{d}$ corresponds to the instability of the phase 1 when the barrier disappears. At $n_{ex}\geq n_{d}$ the phase 1 is metastable; at the even higher pumping $n_{ex}\geq n_{d}$ the energy minimum $W_{1}(\eta)$ drops below the initial $W(0)=0$.
After the instantaneous switching, the excess potential energy transforms into the kinetic energy giving rise to collective pendulum oscillations. After a number of oscillations, the system achieves the minimum corresponding to the pure dynamical transition when $W_{b}(\eta_{d})=0$. The state 1 becomes metastable as the state 2 is well below in energy. Fluctuations of basic parameters result also in formation of alternating domains of coexisting phases. The third critical pumping appears when the spinodal is formed, the barrier merges with the phase 1 minimum which corresponds to the absolute instability of the phase 1. The pumping threshold $n_{c3}$ does not dependent on excitons’ coupling parameters.
The numerical modeling of a similar system can be found in [33].
3 Intrinsic excitons and the excitonic insulator.
3.1 Generic excitonic insulator.
Now we turn to the conceptually more complicated case where the density of excitons and the order parameter $\rho$ are of the same origin; they are additive and statically indistinguishable. This dualism is illustrated by physics of phase transitions from a (quasi)neutral to ionic states in donor-acceptor molecular stacks (see [37] for the latest review and the analysis of the phase diagram, [6, 38] for a review of PIPT experiments, [39] for their later development, and our relevant discussion in [40]). In these compounds, the high-T state already shows a charge alternation $\rho_{1}$ among donor and acceptor sites which is natural because of their different affinities. But at lower T the phase transition takes place to the ”ionic” state characterized by a higher $\rho_{2}>\rho_{1}$ (in units of the electron charge $e$, typically $\rho_{2}\approx 0.7,\rho_{1}\approx 0.3$). Apparently, this transition is not symmetry breaking, but commonly also another, symmetry breaking, order parameter enters the game. This is a spontaneous dimerization of bonds $u$ (akin to (spin)Peierls effect) on top of the either built in or spontaneously augmented dimerization of sites; we shall return to this issue below in Section 4.
The charge transfer present in the ground state can be viewed as a condensate of excitations composed with e-h pairs which otherwise need to be excited across the excitation gap. The appropriate wave function of the e-h state is composed with all possible eigenstates including their free conducting states and bound ones - the Wannier-Mott excitons. But the fortunate fact, that, in the material of the primary interest, the e-h continuum gap lies three times higher than the energy $E_{ex}$ of the excitons, allows us to concentrate only on the later ones. Thus, we arrive at the picture of the Excitonic Insulator suggested long time ago for more conventional semiconductors. The incoherent optical pumping adds a density of excitons which might be additive to the one frozen in to the ground state. But if the added excitons have a time to be cooled down to the BEC state, then both the BE and the EI condensates are described by the unique wave function $\Psi$. This integrity is demonstrated by the duality of theoretical descriptions of the BEC of excitons and of the EI [41]. The formulations differ only by the monitoring parameters: the mean density for the BEC versus the chemical potential for the EI. Microscopically the inerpolation among the BEC and the EI was considered in [42, 43].
Since we concentrate upon a situation with the first order transition, then in equilibrium, without pumping, the energy functional $W(\rho)$ must have two minima: the stable one at $\rho_{1}$ and the metastable one at $\rho_{2}$ separated by the barrier $W_{b}=W(\rho_{b})$ at $\rho_{b}$. The first order phase transition happens at the temperature $T_{c}$ when $W(\rho_{1})=W(\rho_{2})$.  In spite of the essential jump in $\rho$ observed experimentally, there are clear indications that the first order phase transition is close to a second order one. This proximity helpfully allows to use the Landau expansion for the system energy functional which (for a homogeneous state) can be written as
$W(\Psi)$ $=W_{\rho}+W_{phase}(\Psi)$ (6)
$W_{\rho}$ $=E_{ex}^{0}\rho+\frac{a}{2}\rho^{2}+\frac{b}{3}\rho^{3}~{},~{}\rho=\Psi^{\ast}\Psi+n_{incoh}$ (7)
The density $n_{inc}$ of the incoherent cloud of higher energy excitons is taken into account here for further purposes to which we shall return in Section 4. The energy $W_{\rho}$ depends only on the exciton density; its derivative gives the thermodynamic definition of the exciton energy
$\frac{\partial W}{\partial\rho}=U(\rho)=E_{ex}^{0}+a\rho+b\rho^{2}$ (8)
The term $W_{phase}(\Psi)$ is a special energy contribution depending on the phase of $\Psi$ which we shall discuss just below. Notice in (7) the linear dependence of the energy upon the explicit order parameter $\rho$. Actually the energy starts quadratically in therms of the hidden parameter $\Psi$. Also, being small, $\rho$ cannot turn negative because $\rho=\Psi^{\ast}\Psi>0$ which constraint cannot be seen thermodynamically.
Figure 1: Four regions for $W(\rho)$ and $U(\rho)$. I an II: decreasing $U(\rho)>0$, attraction of excitons. II and III: $U(\rho)<0$, spontaneous creation of excitons. III and IV: increasing $U(\rho)>0$, repulsion of excitons
The energy $W(\rho)$ can have two minima: the stable one which we define as $\rho_{1}=0$ and a metastable one at $\rho=\rho_{2}$ separated by the barrier at $\rho_{b}$. This standard form is significant in terms of the excitons as showing four different regimes shown in Fig.1. In the region I $W(\rho)$ grows with $\rho$, hence the exciton energy $E_{ex}(\rho)>0$, but it decreases with increasing $\rho$ which corresponds to the attraction of excitons. In the regions II,III $E_{ex}\eqslantless 0$ indicating the regime of the EI state. Region II and III are separated by the inflation point $\rho^{\ast}$of $W(\rho)$ where $E_{ex}<0$ is minimal; at $\rho>\rho^{\ast}$, the Eex, being still negative, starts to grow which indicates on the excitons’ repulsion. In the region IV we pass the metastable point where $E_{ex}(\rho_{2})=0$ and it grows with increasing of $\rho$ with the feature of the BEC.
Integration of the BEC of excitons and the EI approaches allows for the combined description for the dynamics of excitons and the evolution of the phase transition after the pumping. It include both thermodynamic and dynamic effects of the PIPT. Still, adopting the concept of the EI we arrive at the enigma. The dualism of the excitonic density $|\Psi|^{2}$ and the thermodynamic order parameter $\rho$ do not seem to be consistent. Within the thermodynamic picture, the order parameter $\rho$ is a single real non-conserved field. The expected equation allows for unrestricted evolution towards the energy minimum in $\rho$:
$\frac{d^{2}\rho}{dt^{2}}\varpropto-\frac{\delta W}{\delta\rho}.$
which was very well verified in PIPT science.
The charge transfer density from excitons $n_{ex}=|\Psi|^{2}$ is also the real field which seems to be equivalent or additive to $\rho$. Nevertheless, its evolution is described by a complex wave function $\Psi=|\Psi|\exp(i\zeta)$ of the BEC. The corresponding evolution of the excitonic condensate contains the hidden degree of freedom, the phase $\zeta$. The evolution of $\zeta$ is measurable through the instantaneous energy of the exciton (defined dynamically as $E_{ex}=-\hbar\partial_{t}\zeta$) which in the homogeneous regime (or in average over the space) coincides with the thermodynamic definition $E_{ex}\Rightarrow U={dW}/{dq}$.
The excitons are the interacting bosons which at sufficiently high concentration and low temperature can be described collectively by an adapted Gross–Pitaevskii theory. This is a generalization of the nonlinear Schrödinger equation (NLSE) which conserves the total number of particles, hence no evolution of $\rho$ is allowed.
[…]
To make the evolution of the fields $\Psi$ and $\rho$ compatible, we need to take into account the processes which do not conserve the number of particles, i.e. given by phase-dependent interactions. They appear from anomalous matrix element of Coulomb interactions transferring two electrons across the gap, from filled to empty band [41] and have been studied intensively for bi-layer systems ([44] with a recent return of interests in [45]). That means the simultaneous creation or annihilation of two e-h pairs, i.e. creation/destruction of two excitons from/to the vacuum. Being virtual usually (the energy is not conserved), these transition amplitudes acquire nonzero averages for macroscopic concentrations in the BEC or EI states.
Figure 2: Normal and anomalous Coulomb interactions
Finally, the excitonic insulator free energy acquires an additional contribution which fixes the phase, hence violates the particles number conservation.
$\Delta W_{phase}(\Psi)=\frac{1}{2}(A^{\ast}\Psi^{2}+A\Psi^{\ast 2})=A\rho\cos[2(\zeta-\alpha)]$ (10)
(From here on, we take the material constant $\alpha=0$ with $A>0$ which can always be done by choosing the origin of the variable phase $\zeta$.)
By the definition of the EI, the phase fixing terms are small, $A\ll 1$, hence the total complex order parameter $\Psi$ still needs to be exploited. The relative value of anomalous $A$ terms, with respect to the main Coulomb energy $E_{b}$, is of the order of $A/E_{b}\propto(a/R)^{d}$ where $a$ is the lattice spacing, $R\gg a$ is the exciton radius, and $d$ is the space dimension which is small by definition for Wannier–Mott excitons. For donor-acceptor chains, the bare value of $A$ may not be very small as $R$ is only of a few $a$ and $d=1$. But the fact that the system is nearly one–dimensional results in strong phase fluctuations which reduce functions periodic in $\zeta$. Then the effective value $A\rightarrow A\langle exp(2i\zeta)\rangle\approx A\exp(-2\langle\zeta^{2}\rangle)$ can be small and even be renormalized to zero (in the context of excitons in two-dimensional structures, effects of the spacial phase incoherence where discussed in [11]).
At a low but still macroscopic density of residual excitons, $\rho\rightarrow 0$, $U\rightarrow E_{ex}^{0}$, the energy of a single excitons is shifted down to $E_{A}=\sqrt{(E_{ex}^{0})^{2}-|A|^{2}}$. The wave function oscillates now with two mirror frequencies: ±$\pm\omega_{A}=\pm E_{A}/\hbar$ where the energy $E_{A}<E_{ex}^{0}$ is shifted down with respect to the bare $E_{ex}^{0}$. The eigenfunction of the single exciton appears as the interference:
$\Psi$ $\propto\sqrt{1-|A|/E_{ex}^{0}}\cos(\omega_{A}t)+i\sqrt{1+|A|/E_{ex}^{0}}\sin(\omega_{A}t),$ (11)
$|\Psi|^{2}$ $\propto 1-|A|/E_{ex}^{0}\cos(2\omega_{A}t)$
This quantum interference of positive and negative energies leading to persistent coherent oscillations of the number of excitons was recently interpreted as a dynamical Josephson effect predicted for excitonic condensates in double-well heterostructures or natural bi-layer systems [45, 46].
3.2 Evolution equations for the generic excitonic insulator
We shall adopt the Gross–Pitaevskii type equation which is applicable when, for all relevant states of bosons, their occupation numbers are much bigger than unity. For excitons on a $d$–dimensional lattice, the condition that their mean density per lattice site is small is $n_{ex}\gg(T/D)^{d}$. With the exciton bandwidth $D\sim 10^{2}\ K$, this inequality can always be satisfied for a typical experimental temperatures $T\sim 10^{1}$. Even if the initial value of $n_{ex}$ is not sufficiently high, the kinetics of cooling feeds the low energy states, then the Gross–Pitaevskii theory becomes applicable sooner or later. This is an advantage of the fast PIPT technique, where the time of observations is shorter than the recombination life time of excitons.
The basic equation can be chosen as
$i\hbar\partial_{t}\Psi=-\frac{\hbar^{2}}{2M}\partial_{x}^{2}\Psi+U(|\Psi|^{2})\Psi-A/2\Psi^{\ast}-i\hbar\Gamma\Psi$ (12)
Here appear two terms breaking the conservation of total number of excitons: the nondissipative term $\varpropto A$ and the common dissipative one $\sim\Gamma$. The relaxation rate $\Gamma$ has a sophisticated behavior: depending on $\rho$, $\Gamma$ evolves from a constant $\Gamma=1/\tau_{rec}$ at vanishing $\rho$ (the single-particle recombination of rare excitons) to $\Gamma\propto\rho$ at moderate $\rho$ (Bose–Einstein statistics and also Auger processes, well proved in semiconductors). Nevertheless, around the second minimum of $W$ at $(\rho\approx\rho_{2})$ of the high density equilibrium phase, we expect that $\Gamma(\rho)\rightarrow 0$ since there is no channel for decay. (We neglect the evaporation over the barrier towards the $\rho=0$.) We can interpolate among these regimes writing $\Gamma=G(\rho)\rho U(\rho)/\hbar$, where $G(\rho)$ is some structureless dimensionless function of $\rho$ (which we shall choose as a constant for the numerical modeling). In the region II,III of Fig.1 $\Gamma<0$ corresponding to the creation of excitons - the amplification of the number of excitons instead of their decay. Another way to reason is to notice that the energy relaxation terminates when $\partial_{t}\zeta=0$ which leads to another form of $\Gamma$ equivalent in a space homogeneous regime:
$\Gamma\Rightarrow-\frac{G}{2i}(\Psi^{\ast}\partial_{t}\Psi-\Psi\partial_{t}\Psi^{\ast})=-G\rho\partial_{t}\zeta=\frac{G(\rho)\rho}{\hbar}E_{ex}(t,x),\ G\approx const.$ (13)
3.3 Space homogeneous regime or zero dimension
Enforcing the space homogeneous regime $\nabla\Psi\equiv 0$ or for a a zero-dimensional system, the equations for $\rho,\zeta$ acquire a simple form
$\hbar\partial_{t}\zeta$ $=-U(\rho)+|A|\cos(2\zeta)$ (14)
$\partial_{t}\rho$ $=-\Gamma\rho+2\left|A\right|\rho\sin(2\zeta)=G\rho^{2}\partial_{t}\zeta+\left|A\right|\rho\sin(2\zeta)$ (15)
$=-(G/\hbar)\rho^{2}(U(\rho)-|A|\cos(2\zeta))+2\left|A\right|\rho\sin(2\zeta).$
The static solution $\partial_{t}\Psi=0$ becomes possible taking into account the effect of the phase locking, when
$\sin(2\zeta)=0~{},-U(\rho_{st})+|A|\cos(2\zeta)=0$
i.e.
$\zeta_{st}=\pi n/2\ ,\ U(\rho_{st})=|A|(-1)^{n}$
The presence of the $A$ term results in shifting the static value $\rho_{st}$ with respect to the thermodynamically equilibrium value $\rho_{eq}$.
The dynamics is illustrated by the numerical solution of the above equations. For the modeling we specify the ground state energy as
$W(\rho)=\rho((\rho-1)^{2}+0.05)/1.05\ ,\ U(\rho)=dW/d\rho$
where $W(\rho)$ is normalized to have the bare exciton energy $E_{ex}^{0}=U(0)=1$. With these parameters, we are below the thermodynamic phase transition to the excitonic insulator state but it can exist as a metastable state (the minimum of $W(\rho)$ at $\rho_{2}\approx 1$).
Figure 3 shows time dependencies of the phase and the amplitude, and the trajectory as a parametric polar plot for $\rho(\zeta)$. The sub-barrier pumping leaves the system in the unlocked regime and it relaxes to the virgin ground state with $\rho\rightarrow 0$ and $E_{ex}(t)\rightarrow E_{ex}^{0}$. In the super-barrier regime, $\zeta$ and $\rho$ grow reaching the locking transition where the exciton energy $E_{ex}(t)\rightarrow E_{loc}$ gradually freezes at the thermodynamic equilibrium while the phase is locked via the dynamical transition at an allowed value $\pi(n+1/2)$. For both cases the curves are superimposed by oscillations originated via the $A$ term by destruction/creation of pairs of excitons.
Figure 3: Time evolution of $\rho(t)$, $\zeta(t)$, $\partial_{t}\zeta(t)$ and the polar trajectory $\rho(\zeta)$
3.4 Inhomogeneous regimes and space-time patterns.
Now we open a possibility for a space-inhomogeneous solution considering a one-dimensional system with a length $L$, $-L/2<x<L/2$. The dynamical interplay of various fields results in transient effects like the self–trapping of excitons and the consequent nucleation of regions with increased exciton density. The effect of self-trapping of excitons results in the possibility of local phase transformations already above a threshold $\rho^{\ast}<\rho_{b}$ which is lower than the barrier threshold $\rho_{b}$ for the homogeneous regime. The reason is the collective self–trapping [47] of excitons similar to self–focusing in optics. For $\rho_{ini}<\rho^{\ast}$ the evolution of $\rho$ and $\Psi$ still shows a transient self-trapping, which nevertheless levels out at the long time. For the intermediate initial pumping $\rho^{\ast}<\rho_{ini}<\rho_{b}$, the self-trapping results in a dynamical formation of domains with different states of $\Psi$. Domains grow with time, they experience reflections from the boundaries, collisions and merging. That is accompanied by formation and collapse of domain and domain walls and appearance of propagating wave fronts. The numerical results for a sub-barrier pumping are presented at Fig.4.
Figure 4: Plots for $\rho(t,x)$ at different levels of pumping. A: $\rho_{ini}<\rho_{b}<\rho^{\ast}$. B: $\rho_{b}<\rho_{ini}<\rho^{\ast}$.
4 Excitonic insulator, multi–field model
At appearance of another, symmetry breaking, order parameters, which we associate with lattice displacements $u$, the energy functional acquires additional terms $W(\Psi)\Rightarrow W(\Psi,u)=W(\Psi)+\Delta W(\Psi,u)$ where
$\Delta W(\Psi,u)$ $=\Delta W(\rho,u)+\frac{\beta}{2}\left(\frac{\partial u}{\partial x}\right)^{2}$ (16)
$\Delta W(\rho,u)=\frac{d}{2}(\rho_{u}-\rho)u^{2}+\frac{f}{4}u^{4}$ (17)
Here $\rho_{u}$ is the threshold value of $\rho$ for appearance of displacements $u$. The time evolution of $u$ is governed by the supplementary equation
$K(\partial_{t}^{2}/\omega_{u}^{2}+\gamma\partial_{t})u=-\frac{\delta W(\Psi,u)}{\delta u}$ (18)
Near the BEC-EI crossover, the variable $\Psi$, hence $\rho$, is slow in comparison with $u$ which is governed by the phonon frequency $\omega_{u}$. Then it is helpful to draw the effective potentials $W^{\ast}(\rho)=W(\rho,u_{eq}(\rho))$ and $U^{\ast}(\rho)=U(\rho,u_{eq}(\rho))$ for the equilibrium $u_{eq}(\rho)$ taken after the minimization of $W(\rho,u)$ over $u$ at a given $\rho$. Fig. 5 presents the plots for $W^{\ast}(\rho)$ and $U^{\ast}(\rho)$.
Figure 5: $W^{\ast}(\rho)=W(\rho,u_{eq}(\rho))$ and $U^{\ast}(\rho)=U(\rho,u_{eq}(\rho))$ after the minimization over $u$ at a given $\rho$.
In comparison with the energy $W(\rho)$ for the generic EI in Fig.1, now the energy function $W^{\ast}(\rho)$ demonstrates a new region V at small $\rho<\rho_{c}$, where the exciton energy $E_{ex}$ is positive and it increases with $\rho$ corresponding to the repulsion of excitons. The concave energy curve indicates that there is no self-focusing at small $\rho$, contrary to the situation for generic EI (without lattice deformations), where the convex curve starts from lowest $\rho$ (see Fig.1). The landscape of $W(\rho,u)$ for the full two-field model is presented in Fig.6, to be compared with Fig.6 in ref. [37].
Figure 6: Density plot of $W(\rho,u)$. Left: $\rho_{u}<0$, Right: $\rho_{u}>0$.
At $ho_{u}<0$ only the trivial minimum $\rho=0,u=0$ does exist. At $\rho_{u}>0$ the double metastable minimum appear at $u=\pm u_{0}=\pm\sqrt{(\rho_{u}-\rho)d/f}$. At a higher $\rho>\rho_{c}$ the double minima at $(\rho,$$\pm u_{0})$ become stable while the trivial one $(\rho=0,u=0)$ becomes metastable.
4.1 Time evolution from the numerical modeling.
For the combined model with two order parameters we shall present results of modeling for only homogeneous regimes. For the subcritical pumping $\rho_{ini}<\rho_{u}<\rho_{2}$ with the unperturbed initial $u_{ini}=0$, both $\rho$ and $u$ vain at large $t$. Small oscillations of the exciton energy appear from the macroscopic quantum interference: transfers of pairs of excitons from/to the vacuum. During all the time, $\rho(t)$ decreases monotonically while $E_{ex}$ stays nearly constant close to the unperturbed value. For the supercritical pumping $\rho_{ini}>\rho_{u}$, $\rho$ rises above the initial pumping level by generation of excitons. With $A=0.01$ we observe strong unharmonic oscillations at intermediate times. Sharp lock-in transition gives rise to the new set of strong oscillations. The corresponding time dependencies are shown at Fig.7.
Figure 7: $\rho(t)$ (blue), $u(t)$ (green) $E_{ex}(t)$ (red) for a subcritical (left panel) and a supercritical (right panel pumping.
5 Effects of hot excitons
Above, we have assumed that a quasi–condensate of optically pumped excitons appears sufficiently early as a macroscopic quantum state. It is desirable to also take into account the normal, non condensed density of excitons and its (re)conversion (from)to the condensate. Indeed, temperatures in the experiments are comparable with the estimated degeneracy temperature of the BEC, and they are even higher just after the pumping pulse. The not–quite–resonance pumping also contributes to the initial incoherent density. Within this phenomenologically oriented article, we stay away from the very important question of kinetics of non-condensed particles, relying on experimental facts of fast initial equilibration typical for PIPT.
A future microscopical study can be advanced thanks to the progress in theory of equilibration in a gas of excitons and polaritons [48, 49, 16], and to the general understanding of a non equilibrium Bose gas motivated by problems in cold atoms [14, 15]. For the ideal model of the weakly interacting Bose gas, there is a fair overlap between the regime of the microscopical kinetic and the collective NLSE–based descriptions. But the price is that turbulent mixing must be taken into account in the NLSE [13, 50] and the Gross–Pitaevskii equation must be considered stochastically rather than deterministically [16]. For applications in solid state physics, the universal Bose gas mode is not much helpful as regards the BEC of excitons, because other channels of the relaxation like the emission of phonons [51] become more important than collisions of bosons.
To elucidate these effects we shall present below two types of modeling: i. stochastic equations taking into account the particles falling to the condensate as a random noise; ii. phenomenological description for a dynamical equilibrium with explicit exchange among the condensate and the incoherent condensate.
5.1 Stochastic modeling at $d=0$.
The conversion of particles from the normal to condensate reservoirs and the related decoherence can be modeled numerically for a zero-dimensional system. We need to generalize the NLSE to a stochastic Wigner- Stratonovich process. The type of the random process must be chosen to correspond to random acts of absorption of non-condensed particle into the condensate. The random function $\nu$ needs to be added to the equation for the condensate density $\rho$. That does not affect directly the equation for the phase, but the final effect of the randomness upon the phase will be the most profound as it is demonstrated by the numerical solution presented below.
In calculations, we used the Ito process:
$d\rho=(-G\rho^{2}d\zeta+A\rho\sin(2\zeta))dt-d\nu~{},~{}d\zeta=(-U(\rho)+A\cos(2\zeta))dt$ (19)
The random function $\nu$ describes in average the exponential exhaustion of the reservoir of the normal particles, emulating their diffusion in energy space down to the condensate.
Fig.9 presents the plots for the phase and the density of the condensate and for the corresponding realization of random process (chosen as the exponential Brownian motion). The random increments of $\rho$ are transferred to phase kicks, hence to the decoherence. Another effect, seen at higher $T$ is the sequence of lock-in attempts with switchings among different lock-in values of the phase.
Figure 8: Plots of $\rho(t),$ $u(t)$,$E_{ex}(t)$ for a subcritical (left panel) and a supercritical (right panel) pumping.
5.2 The account for incoherent excitons and a two–fluid description
Hidden problem in experimental feasibility of the BEC-EI scenario is the transformation of the initially incoherent ensemble of pumped excitons to their coherent Bose condensate. Initially hot excitons are expected to form a conventional static ensemble of Bose particles, which occupation numbers of normal states $|k>$ are $n_{k}\sim 1$ with respect to the total large number of particles; their wave functions $\psi_{k}\varpropto\exp(ikx-E_{k}t)$. Being allowed to be quasi-static, which requires for some bottlenecks, this ensemble can be described by Boltzmann kinetics. A more complicated quasi-stationary version is a diffusive energy flow towards low energies. Under this cooling process an intermediate state of a quasi-condensate is formed where at lowest $E_{k}$ the occupation numbers start to be large $n_{k}>>1$. This part of the ensemble can be described by a quasi-classical Gross-Pitaevskii equation, still in the turbulent regime for the condensate with the kinetics for the down-flow of normal particles. At the final equilibration the occupation numbers of lowest eigenstates $n_{0}\sim N$ and they can be described by the single wave function augmented by the equilibrium Bose-Einstein distribution (with the chemical potential locked at $E_{0}$) for normal particles.
The theory is simplified by considering the normal component as a separate quasi–equilibrium reservoir, which can be characterized by the density $n$ or the chemical potential $\mu_{n}$. This approach is validated by particular spectra of polaritons, but it still needs to be justified for our system; we briefly outline its possible application below as an absolutely minimalistic description. A certain ground for the separation into two distinct, particle–exchanging reservoirs comes from suggesting a bottleneck — a minimum $E_{min}$ of the kinetic energy — where the pumped excitations are accumulated after the initial rapid cooling. It is tempting to associate $E_{min}$ with the energy of the lowest lattice mode interacting with excitons. In our case, a good candidate is the soft mode in the dip of the Kohn anomaly, which should exist as a precursor for the lattice dimerization instability, (see [52] and the references therein). That can also be the Debye frequency of the acoustical spectrum; both candidates converge to $E_{min}\sim 100\ K$.
We can make a simplifying, and quite plausible, suggestion that all reservoirs of excitons contribute additively to the order parameter: the total charge transfer becomes $\rho_{tot}=\rho+n$, where still $\rho=|\Psi|^{2}$. Then the system energy and the particle potential become simply $W(\rho+n)$ and $U(\rho+n)$ adding $\mu_{n}=U(|\Psi|^{2}+n)+E_{min}$. Now the equation (12) is further generalized to
$i\hbar\partial_{t}\Psi=-\frac{\hbar^{2}}{2M}\partial_{x}^{2}\Psi+U(|\Psi|^{2}+n)\Psi+\frac{i}{2}Rn\Psi-A/2\Psi^{\ast}$
where $R$ is a conversion rate regulating the exchange among the reservoirs. This equation needs to be complemented by an equation for $n$, which we choose as a simple rate equation with diffusion (cf. [48])
$\partial_{t}n-\partial_{x}b\partial_{x}\mu_{n}n=I-Rn|\Psi|^{2}~{},~{}|\Psi|^{2}=\rho$
[…]
$\mu_{c}=-\hbar\partial_{t}\zeta+\frac{\hbar^{2}}{2M}(\partial_{x}\zeta)^{2}~{},~{}\mu_{n}=U(|\Psi|^{2}+n)+E_{min}~{},~{}R=kn(\mu_{n}-\mu_{c})/\hbar,$
The simpler space–independent Eqs. (14,15) are generalized as
$\hbar\dot{\zeta}$ $=-U(\rho+n)+A\cos(2\zeta),$ (20)
$\dot{\rho}$ $=\rho R+A\rho\sin(2\zeta)=k\rho n(\dot{\zeta}+\mu_{n}/\hbar)+A\rho\sin(2\zeta),$
$=(k/\hbar)\rho n(E_{min}+A\cos(2\zeta))+A\rho\sin(2\zeta)$ (21)
$\dot{n}$ $=I-k\rho n(\dot{\zeta}+\mu_{n}/\hbar)=I-(k/\hbar)\rho n(E_{min}+A\cos(2\zeta)).$ (22)
The results for the case of pumping to the reservoir of uncondensed excitons are presented in Fig. LABEL:fig:9,_fig.10.
Figure 9: Subcritilal pumping, $n(0)=0.27$. Left panel : Plots for $\rho(t)$, $n(t)$, $E_{ex}(t)$,$\pi\zeta(t)/100$; Right panel: Parametric polar plot $(\rho,\zeta)$
Figure 10: Supercritical pumping, $n(0)=0.28$. Left panel: Plots for $\rho(t)$, $n(t)$, $E_{ex}(t)$,$\pi\zeta/2(t)$ Right panel: Parametric polar plot $(\rho,\zeta)$.
At a subcritical pumping $n(0)=0.27$ all quantities show oscillations which are particularly pronounced in $E_{ex}(t)$. The instantaneous frequency of oscillations grows as $f(t)/\pi$. The phase decreases monotonically, correspondingly $E_{ex}(t)>0$, hence the regime of the BEC. The number of condensed particles $\rho(t)$ initially grows, being fed by the decreasing population $n(t)$ of the normal reservoir, then it passes through the maximum and vanishes at long time where $E_{ex}(t)\rightarrow E_{ex}^{0}(t)$. The polar plot shows only one trivial circulation around $\rho=0$.
For supercritical pumping $n(0)=0.28$ the phase start to decrease correspondingly $E_{ex}(t)>0$ for this initial regime of the BEC. Passing through the minimum, the phase starts to grow, correspondingly $E_{ex}(t)<0$ indicating the regime, still nonstationary, of the EI. Finally the phase is locked at the value $\zeta(t)\rightarrow\pi$. The number of condensed particles $\rho(t)$ shows an average growth superimposed by regular oscillations; it keeps growing even after the population $n(t)$ of the normal reservoir is exhausted; later the growth saturates at the equilibrium position $\rho_{st}\approx 1$ of the static EI state. $E_{ex}(t)$ shows the most irregular behavior to finally saturate at $E_{ex}(t)\rightarrow 0$ as it should be for the static EI state. The polar plot shows two circulations: first around $\rho=0$, then switching to the one around $\rho=1$.
For both cases we have presented the results very close to the dynamical pumping threshold which happens sharply between $n_{0}=0.27$ and $n_{0}=0.28$ well below the thermodynamical threshold at $q_{0}=0.36$.
6 Conclusions
The concepts and methods of pump-induced phase transitions can be applied to non-metallic systems where ubiquitous optical excitons can play a role of the pumping agents, unlike the most common practice of pumping to unbound e-h pairs. If the phase diagram of studied material shows a multitude of stable and metastable states then the pumping can achieve switching among these states or even create new static or stationary ones. In general, the excitons and the thermodynamical order parameter can be independent while mutually affecting as we have briefly outlined in Ch.2. But mostly we payed attention to a more specific class which is of a particular interest for theory and of strong experimental implications. Experimentally, this is the case of neutral-ionic transitions in donor-acceptor structures which allow for transformations by both thermodynamical means and the optical pumping. Specifically, here the excitation and the long range ordering are built from the same intermolecular electronic transfers, so that the density of pumped excitons contributes additively to the thermodynamic order. Both thermodynamical and dynamical effects can be described on the same root by viewing the ordered state as the Excitonic Insulator which appears as a macroscopic quantum state of excitons starting as their Bose condensate allowed at high pumping densities. The double nature of the ensemble of excitons leads to an intricate time evolution: the dynamical transition between number–preserved and phase–locked regimes, macroscopic quantum oscillations from interference between the Bose condensate of excitons and the ground state of the excitonic insulator. Modelling of an extended sample shows also stratification in domains of low and high densities which evolve through local dynamical phase transitions and a sequence of domains’ merging. The locally enhanced density of excitons can surpass a critical value to trigger the local persisting phase transformation, even if the mean density is below the required threshold. Beyond our phenomenological descriptions, there are challenging questions the relaxations of initial hot reservoir of excitons to the growing condensate and the establishing of the coherence. The microscopic aspects of these difficult questions have been completely resolved even in much simpler systems like cold atoms or polaritons, Here we have suggested two kinds of illustrations. One is the stochastic generalization of the macroscopic equations to emulate random absorption of incoherent excitons to the condensate. Another illustration is rate equations regulating exchange of normal and condensed reservoirs controlled by a mismatch of chemical potentials. We believe that the presented study and discussions will attract attention to still ill explored excitonic channels in the PIPT science.